solve the following equation..
0 0-1
3 x 0 =x^2
2 0 3

Answer by fractalier(6550)

This is tricky to do here, but if we evaluate the matrix on the left, we get
(-1)(0-2x)=x^2
2x = x^2
x = 0 or x = 2